OpenStudy (anonymous):

P = 2l + 2w. Solve for l. Awarding medal

OpenStudy (squirrels):

So what you want to do is isolate "l". Having it on one side of the equation.

OpenStudy (anonymous):

Can u write it out for me?

OpenStudy (squirrels):

To do that, lets subtract 2w from both sides. \[P-2w=2l\]

OpenStudy (squirrels):

Now to get "l" by itself, divide "2" from both sides. Can you try that?

OpenStudy (anonymous):

p/2 - 2w = 1?

OpenStudy (anonymous):

@Squirrels

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

Hey @Legends is that the answer?

TheSmartOne (thesmartone):

Sorry.

TheSmartOne (thesmartone):

We don't ust give you answers.

TheSmartOne (thesmartone):

OpenStudy is all about the "learning process" you won't learn anything from an answer.

OpenStudy (anonymous):

I asked if that was the right answer. I came up with it myself.

TheSmartOne (thesmartone):

Okie. Let me try to help you.

OpenStudy (anonymous):

Ok. Thanks

TheSmartOne (thesmartone):

So we have \(\sf P = 2l + 2w\)

OpenStudy (anonymous):

and we're solving for l

TheSmartOne (thesmartone):

To get l all by itself. We should first subtract 2w on both sides. What will we get if we subtract 2w from both sides?

OpenStudy (anonymous):

p-2w = 2l?

OpenStudy (legends):

sorry @purpleowl i wasn't paying attention

OpenStudy (anonymous):

its ok. @TheSmartOne is that right?

TheSmartOne (thesmartone):

Yes that is :)

TheSmartOne (thesmartone):

So now we need to divide both sides by 2. So currently we have \(\sf p-2w = 2l\)

OpenStudy (anonymous):

is the answer p - 2w/2 = l?

OpenStudy (anonymous):

@TheSmartOne are you there??

OpenStudy (legends):

It looks like you got it

TheSmartOne (thesmartone):

Nope

TheSmartOne (thesmartone):

Its not correct.

TheSmartOne (thesmartone):

\(\sf p-2w = 2l\) Since we have this, we have to divide 2 on both sides. \(\sf\Large\frac{p-2w}{2} = \frac{2l}{2}\) And if we simplify this we get \(\sf\Large\frac{p}{2}-\frac{2w}{2}= l\)

TheSmartOne (thesmartone):

So we can't simplify \(\sf\Large\frac{p}{2}\) anymore. But we can simplify \(\sf\Large\frac{2w}{2}\). Notice the 2 in the numerator and the denominator? We can cross them out and simplify it down to \(\sf\large w\)

TheSmartOne (thesmartone):

So at the end we have \(\Large\sf l = \frac{P}{2} - w\)

TheSmartOne (thesmartone):

@purpleowl That help you? :)